OpenStudy (anonymous):

The sum of two numbers is 20. If one number is x, then the other number is ?. their product is p(x)=?. find the maximum value of p.

OpenStudy (anonymous):

x + y = 20 The product is (x)(20 - x) = p(x) You can now get the maximum either through the vertex form for a parabola opening downward or you can get it by differential calculus.

OpenStudy (anonymous):

Can you perform either method?

OpenStudy (anonymous):

Yup, I got it now. I was just unsure of how to set it up! Thanks :D

OpenStudy (anonymous):

you're quite welcome!